OpenStudy (anonymous):

y-1=2(x-2) write in standard form

OpenStudy (across):

y-1=2(x-2), y-1=2x-4, y=2x-3. Easy enough?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

y-1=2(x-2)is in Point slope form. To get it into standard form get rid of the () by distributing the 2 and move the term 2x to the left and the constants to the right y - 1 = 2 * x + 2(-2) -2x + y = -4 + 1 -2x + y = -3 = Standard Form